# Connectivity-Preserving Multi-Agent Area Coverage via Optimal-Transport-Based Density-Driven Optimal Control (D²OC)

### v1

@@ I Introduction
Density-Driven Optimal Control (D2OC): This work builds on the D2OC framework [12, 13], where a reference density defines spatial priorities and the Wasserstein distance measures how well the agent distribution aligns with this reference. Control inputs steer agents to reduce this discrepancy while respecting dynamics and motion constraints. Unlike [12, 13], which derived an optimal-control solution using Lagrange multipliers, the present work reformulates the problem to enable a quadratic-program (QP) representation and further incorporates connectivity-preserving constraints to maintain inter-agent communication during coverage.
Contribution: This work enhances D2OC by enforcing connectivity during non-uniform coverage. The main contributions are summarized as follows: 1) QP Equivalence and Convexity Analysis: The Wasserstein-based D2OC objective, without connectivity constraints, is shown to be equivalent to a QP, and its strict convexity and associated optimal solution, including the closed-form unconstrained optimizer, are established. These results were not presented in [12, 13]; 2) Connectivity-Preserving Penalty with Reachable Sets: A smooth connectivity penalty integrated with a reachable-set formulation is developed to maintain inter-agent communication without enforcing rigid formations. When this penalty is included, the overall formulation remains convex but is no longer quadratic, an aspect absent from [12, 13]; 3) Simulation-Based Validation: Representative simulations confirm that the proposed connectivity-preserving D2OC achieves improved coverage efficiency, faster convergence, and sustained communication compared with the unconstrained case.
@@ II-A Wasserstein Distance and Optimal Transport
subject to the constraints $\pi_{\ell j}\geq 0,\,\sum_{j=1}^{N}\pi_{\ell j}=\alpha_{\ell},\,\sum_{\ell=1}^{M}\pi_{\ell j}=\beta_{j},\,\sum_{\ell,j}\pi_{\ell j}=1,$ where $\pi_{\ell j}$ specifies the amount of probability mass transported from $y_{\ell}$ to $q_{j}$. In this paper we focus on the quadratic case ($p=2$) with Euclidean distance as the ground cost for $d(\cdot)$.
@@
$\rho(k)=\tfrac{1}{k+1}\sum_{t=0}^{k}\bigg(\tfrac{1}{n_{a}}\sum_{i=1}^{n_{a}}\delta_{y_{i}(t)}\bigg),\,\nu=\tfrac{1}{N}\sum_{j=1}^{N}\delta_{q_{j}},$ (3)
@@ II-B Decentralized Coverage Protocol
The first two stages are performed independently by each agent, while the third stage coordinates neighboring agents via weight sharing. Iterating this cycle gradually aligns the agents’ empirical distribution with the reference distribution. In this work, we focus on deriving the optimal control input for the first stage under a connectivity-preserving constraint. A full description of the overall cycle is provided in [4].
@@ III Formulation of the D2OC Cost Function via Quadratic Programming
This defines the number of time steps required for the control input $u_{i}(k)$ to first have a direct effect on the output vector $y_{i}(k)$. Then, the cost function for agent $i$ using the squared local Wasserstein distance to achieve D2OC is defined over the prediction horizon $H\in\mathbb{Z}_{>0}$, starting from time $k+r$:
@@ IV-A Connectivity Constraint with Reachable Sets
$\mathcal{Z}^{y}_{j}(k+h)=\hat{y}_{j}(k+h)\oplus G_{j}(k+h)\mathcal{B}_{\infty},$ (13)
where $\hat{y}_{j}(k+h)$ denotes the nominal prediction of agent $j$’s output at time $k+h$, obtained from its known model $(A_{j},B_{j},C_{j})$ and the most recently exchanged output. This prediction is used as the reference center of the reachable set for enforcing the connectivity constraint. The generator matrix $G_{j}(k{+}h)\!\in\!\mathbb{R}^{d\times m}$ represents the linear mapping from bounded control deviations to predicted output deviations, computed as
@@
which applies for all $h\geq r$, as inputs affect the output only from step $r$ onward. For $h<r$, $G_{j}(k{+}h)$ is the zero matrix.
Each column of $G_{j}(k{+}h)$ represents the effect of one control-input direction propagated through the dynamics, and $\mathcal{B}_{\infty}=\{z\in\mathbb{R}^{m}:\|z\|_{\infty}\leq 1\}$ is the unit hypercube capturing all admissible combinations of these bounded deviations. Equivalently, $\mathcal{Z}^{y}_{j}(k+h)=\{\hat{y}_{j}(k+h)+G_{j}(k+h)z\mid\|z\|_{\infty}\leq 1\}.$
@@
$R_{j}(k+h)=\textstyle\max_{\|z\|_{\infty}\leq 1}\|G_{j}(k+h)z\|,$
and enforce $\|y_{i}(k+h)-\hat{y}_{j}(k+h)\|\leq r_{\mathrm{comm}}-R_{j}(k+h)$, $h=r,\dots,r+H-1$, which ensures the reachable set of agent $j$ lies within agent $i$’s communication ball while accounting for bounded inputs. To enable fast online connectivity checks without handling the full zonotope geometry, we adopt a conservative scalar radius bound:
$R_{j}(k{+}h)=\max_{\ell}\|G_{j,\ell}(k{+}h)\|_{2},$ (15)
where $G_{j,\ell}(k{+}h)$ denotes the $\ell$-th column of $G_{j}(k{+}h)$.
@@ IV-B Soft-Constraint for Connectivity-Preserving Control
We relax the feasibility requirement by introducing a soft penalty on communication range violations. Specifically, for each horizon step $h\in\{r,\dots,r+H-1\}$, we define the predicted output
@@
$J_{\mathrm{soft}}(U_{i}^{k|H})=$ $\frac{1}{2}(U_{i}^{k|H})^{\top}HU_{i}^{k|H}+f^{\top}U_{i}^{k|H}+\sum_{h=r}^{r+H-1}\phi(T_{i}(h)).$ (22)
@@
The quadratic term $\frac{1}{2}(U_{i}^{k|H})^{\top}HU_{i}^{k|H}$ is strictly convex. Adding convex penalties preserves strict convexity. Restricting a strictly convex function to the convex set $\mathcal{U}_{i}^{k|H}$ yields a unique minimizer.
@@
The soft-constraint formulation provides a convex relaxation of the hard connectivity condition, allowing limited violations while maintaining overall communication. The penalty parameters $\kappa$ and $\eta$ jointly determine the strength and sharpness of this relaxation. Increasing $\kappa$ reinforces the penalty and improves connectivity preservation, whereas larger $\eta$ produces a steeper transition toward a hard constraint and may introduce numerical stiffness if too large. In practice, $\kappa$ is increased until violations are consistently penalized, and $\eta$ is then selected as the smallest value that sharpens the penalty without causing oscillatory or saturated behavior. Smaller $\eta$ values yield smoother but weaker enforcement, whereas larger $\eta$ values impose stricter connectivity at the cost of reduced smoothness.
@@ V Simulations
(a) Trajectories without connectivity constraint
(b) Inter-agent distances without connectivity constraint
@@
Without the connectivity constraint (Figs. 1a,b), agents moved freely and dispersed widely across the domain. Although the communication range was $r_{\mathrm{comm}}=15$, the absence of constraint allowed inter-agent distances to exceed 200, as shown in Fig. 1(b). Trajectories appeared to cover the reference samples (green) uniformly. However, since connectivity was not enforced, agents could share weight information, which represents local coverage, only when they temporarily came within range of others, resulting in event-based communication. Once disconnected, they lost awareness of covered regions, causing redundant revisits and inefficient exploration.
@@ VI Conclusion
Future work includes extending the framework to dynamic environments with time-varying communication and sensing conditions and incorporating emergency protocols for restoring connectivity when links are lost. We also plan to conduct experimental validation of the proposed connectivity-preserving D2OC framework using a multi-agent platform.
Acknowledgements
This work was supported by NSF CAREER Grant CMMI-DCSD-2145810.

### v3

@@ I Introduction
Density-Driven Optimal Control (D2OC): This work builds on the D2OC framework [12], where a reference density defines spatial priorities and the Wasserstein distance measures how well the agent distribution aligns with this reference. Control inputs steer agents to reduce this discrepancy while respecting dynamics and motion constraints. Unlike [12], which derived an optimal-control solution using Lagrange multipliers, the present work reformulates the problem to enable a quadratic-program (QP) representation and further incorporates connectivity-preserving constraints to maintain inter-agent communication during coverage.
Contribution: This work enhances D2OC by enforcing connectivity during non-uniform coverage. The main contributions are summarized as follows: 1) QP Equivalence and Convexity Analysis: The Wasserstein-based D2OC objective, without connectivity constraints, is shown to be equivalent to a QP, and its strict convexity and associated optimal solution, including the closed-form unconstrained optimizer, are established. These results were not presented in [12]; 2) Connectivity-Preserving Penalty with Reachable Sets: A smooth connectivity penalty integrated with a reachable-set formulation is developed to maintain inter-agent communication without enforcing rigid formations. When this penalty is included, the overall formulation remains convex but is no longer quadratic, an aspect absent from [12]; 3) Simulation-Based Validation: Representative simulations confirm that the proposed connectivity-preserving D2OC achieves improved coverage efficiency, faster convergence, and sustained communication compared with the unconstrained case.
@@ II-A Wasserstein Distance and Optimal Transport
subject to the constraints
$\pi_{\ell j}\geq 0,\,\sum_{j=1}^{N}\pi_{\ell j}=\alpha_{\ell},\,\sum_{\ell=1}^{M}\pi_{\ell j}=\beta_{j},\,\sum_{\ell,j}\pi_{\ell j}=1,$
where $\pi_{\ell j}$ specifies the amount of probability mass transported from $y_{\ell}$ to $q_{j}$. In this paper we focus on the quadratic case ($p=2$) with Euclidean distance as the ground cost for $d(\cdot)$.
@@
$\rho(k)=\tfrac{1}{k+1}\sum_{t=0}^{k}\bigg(\tfrac{1}{n_{a}}\sum_{i=1}^{n_{a}}\delta_{y_{i}(t)}\bigg),\,\nu=\frac{1}{N}\sum_{j=1}^{N}\delta_{q_{j}},$ (3)
@@ II-B Decentralized Coverage Protocol
The first two stages are executed independently by each agent (decentralized control), while the third stage enables coordination through weight sharing among neighbors. Repeated execution of this cycle gradually aligns the empirical distribution of agents with the reference distribution over their operational horizon. The objective of this study is to determine the optimal control input under a connectivity-preserving constraint in the first stage, while the full description of each stage can be found in [12].
@@ III Formulation of the D2OC Cost Function via Quadratic Programming
This defines the number of time steps required for the control input $u_{i}(k)$ to first have a direct effect on the output vector $y_{i}(k)$. Then, the cost function for agent $i$ using the squared local Wasserstein distance to achieve D2OC is defined over the prediction horizon $H\in\mathbb{Z}$, starting from time $k+r$:
@@ IV-A Connectivity Constraint with Reachable Sets
$\mathcal{Z}^{y}_{j}(k+h)=\hat{y}_{j}(k+h)\oplus G_{j}(k+h)\,\mathrm{diag}(\Delta u_{j})\,\mathcal{B}_{\infty}$, (13)
where $\hat{y}_{j}(k+h)$ denotes the nominal prediction based on $(A_{j},B_{j},C_{j})$ and the most recently exchanged output. Here, $\mathcal{B}_{\infty}=\{z\in\mathbb{R}^{m}:\|z\|_{\infty}\leq 1\}$ is the unit hypercube scaled by the half-range $\Delta u_{j}=(u_{\max}-u_{\min})/2$, where $u_{\max}$ and $u_{\min}$ are the maximum and minimum admissible inputs. Note that the input shift $\hat{u}_{j}=(u_{\max}+u_{\min})/2$ is not included in (13), which is required for the exact reachable set, while it is unnecessary for computing the conservative radius of the reachable set. The generator matrix $G_{j}(k{+}h)\in\mathbb{R}^{d\times m}$ maps bounded control deviations to output deviations:
@@
with $G_{j}(k{+}h)=0$ for $h<r$. Each column represents the effect of one input direction.
Equivalently,
$\mathcal{Z}^{y}_{j}(k+h)=\{\hat{y}_{j}(k+h)+G_{j}(k+h)\,\mathrm{diag}(\Delta u_{j})\,z\mid\|z\|_{\infty}\leq 1\}.$
@@
$R_{j}(k+h)=\max_{\|z\|_{\infty}\leq 1}\|G_{j}(k+h)\,\mathrm{diag}(\Delta u_{j})\,z\|,$
and enforce $\|y_{i}(k+h)-\hat{y}_{j}(k+h)\|\leq r_{\mathrm{comm}}-R_{j}(k+h)$, $h=r,\dots,r+H-1$, ensuring that the reachable set lies within agent $i$’s communication ball.
For fast online checks, a conservative bound is
$R_{j}(k{+}h)=\max_{\ell}\|G_{j,\ell}(k{+}h)\,\Delta u_{j}\|,$
where $G_{j,\ell}(k{+}h)$ is the $\ell$-th column of $G_{j}(k+h)$.
@@ IV-B Soft-Constraint for Connectivity-Preserving Control
We relax the feasibility requirement by introducing a soft penalty on communication range violations. Specifically, for each horizon step $h\in\{r,\dots,r+H-1\}$, let the predicted output be
@@
$J_{\mathrm{soft}}(U_{i}^{k|H})=$ $\frac{1}{2}(U_{i}^{k|H})^{\top}H_{i}U_{i}^{k|H}+f_{i}^{\top}U_{i}^{k|H}+\sum_{h=r}^{r+H-1}\phi(T_{i}(h)).$ (21)
@@
The quadratic term $\frac{1}{2}(U_{i}^{k|H})^{\top}H_{i}U_{i}^{k|H}$ is strictly convex. Adding convex penalties preserves strict convexity. Restricting a strictly convex function to the convex set $\mathcal{U}_{i}^{k|H}$ yields a unique minimizer.
@@
The soft-constraint formulation provides a convex relaxation of the hard connectivity condition, allowing limited violations while maintaining overall communication. The penalty parameters $\kappa$ and $\eta$ jointly determine the strength and sharpness of this relaxation. Increasing $\kappa$ reinforces the penalty and improves connectivity preservation, whereas larger $\eta$ produces a steeper transition toward a hard constraint and may introduce numerical stiffness if too large. In practice, $\kappa$ is increased until violations are consistently penalized, and $\eta$ is then selected as the smallest value that sharpens the penalty without causing oscillatory or saturated behavior. Smaller $\eta$ values yield smoother but weaker enforcement, whereas larger $\eta$ values impose stricter connectivity at the cost of reduced smoothness. A margin $\gamma\in(0,1)$ scales the communication range $r_{\mathrm{comm}}$ to reduce soft-constraint violations.
@@ V Simulations
(a) Trajectories without constraint
(b) Inter-agent distances without constraint
@@
Without the connectivity constraint (Figs. 1a,b), agents moved freely and dispersed widely across the domain. Although the communication range was $r_{\mathrm{comm}}=15$, the absence of constraint allowed inter-agent distances to reach as high as 200, as shown in Fig. 1(b). Trajectories appeared to cover the reference samples (green) uniformly. However, since connectivity was not enforced, agents could share weight information, which represents local coverage, only when they temporarily came within range of others, resulting in event-based communication. Once disconnected, they lost awareness of covered regions, causing redundant revisits and inefficient exploration.
@@ VI Conclusion
Future work includes extending the framework to dynamic environments with time-varying communication and sensing conditions, and incorporating emergency protocols for restoring connectivity when links are lost. We also plan to conduct experimental validation of the proposed connectivity-preserving D2OC framework using a multi-agent platform.